# Characterizing Density and Gravitational Potential Fluctuations of the Interstellar Medium

### v1

@@ 1 Introduction
However, in reality, individual ISM substructures are observed to be far from spherically symmetric, often instead taking a more extended, filamentary form (e.g., J. D. Soler et al. 2022; A. Hacar et al. 2023 and references therein), and substantial fluctuations are observed in the Galaxy across a wide range of scales in atomic as well as molecular gas, and also in dust (e.g., Planck Collaboration et al., 2015). As resolution has improved, surveys of external galaxies in both the Local Group and to distances of tens of Mpc have also revealed extensive structure in both the gas and the dust (e.g., A. D. Bolatto et al., 2008; E. Rosolowsky et al., 2021; J. C. Lee et al., 2023). Thus, empirical characterization of ISM fluctuations now often take the form of multi-scale statistics rather than GMC (or clump) mass functions (see Section 5.2).
In parallel with the development of higher resolution observations that enable more sophisticated empirical characterization of ISM structure, the advent of numerical hydrodynamics as a tool to investigate the ISM has led to vastly more realistic theoretical models. Different kinds of simulations provide different kinds of insight to ISM structure. Global galaxy and cosmological zoom simulations are able to capture important large-scale, long-lived features such as grand design spiral arms by solving N-body and fluid dynamics equations together, and including tidal gravity from nearby companion systems. However, these simulations generally lack the spatial resolution and key ISM microphysical processes that would be needed to realistically model ISM fluctuations at scales below $\sim 100$ pc (with $\lesssim 10$ Myr lifetimes), since the computational cost would be prohibitive for galaxies larger than dwarfs (e.g. R. S. Somerville & R. Davé, 2015; T. Naab & J. P. Ostriker, 2017).
@@
In this work, we characterize ISM fluctuations that develop in the state-of-the-art TIGRESS-NCR simulations (C.-G. Kim et al., 2023). These simulations incorporate the effects of galactic rotation and magnetic fields while employing self-consistent star formation and feedback (ray-tracing radiation and supernovae) as well as nonequilibrium chemistry and cooling, with a simulation resolution of $2-4$ pc. The simulations we analyze span both solar neighborhood and inner-galaxy conditions relevant in a variety of dynamical contexts. With an eye toward extragalactic observations of face-on disk systems (as in e.g., the PHANGS project, A. K. Leroy et al. 2021; J. C. Lee et al. 2023), we build a model for the ISM’s 2D surface density fluctuations, characterizing both their one-point and two-point statistics via probability density functions (pdfs) and power spectra. Additionally, we characterize the time dependence at each spatial frequency. For use in stellar dynamical applications, we also characterize the ISM’s 3D gravitational potential fluctuations.
The rest of this paper is organized as follows. First, in Section 2, we describe in greater detail the TIGRESS-NCR simulations that we study throughout this work. In Section 3 and Section 4 respectively, we characterize the ISM’s surface density fluctuations and connect them to the structure of the ISM’s volume density and gravitational potential. We discuss interpretations, implications, and caveats of our results, and connect with observations in Section 5. Finally, we summarize in Section 6.
2 Simulations
@@
The TIGRESS simulations employ the “shearing box” approximation (P. Goldreich & D. Lynden-Bell, 1965; J. F. Hawley et al., 1995) to model a locally Cartesian patch of the Galactic disk centered at a Galactocentric radius $R_{0}$, rotating at the local circular orbital frequency $\Omega$. The in-plane coordinates $x\equiv R-R_{0}$ and $y\equiv R_{0}(\varphi-\Omega t)$ represent the radial and azimuthal directions respectively, while the $z$ coordinate represents the distance above or below the midplane. The simulation domain extends a length $L_{x}=L_{y}\equiv L$ in each of the in-plane directions and $L_{z}$ in the vertical direction, with a cell size of $\Delta x$ in all dimensions, so that there are $N_{x}=N_{y}=L/\Delta x$ cells in each planar direction, and $N_{z}=L_{z}/\Delta x$ cells resolving the vertical direction. The background differential rotation of the galaxy is modeled as a linear shear flow, so that a circular orbit at a radial position $x$ has a velocity $\mathbf{v}_{0}=-q\Omega x\hat{y}$, where $q\equiv-(\mathrm{d}\ln\Omega/\mathrm{d}\ln R)_{R_{0}}$ parametrizes the shear. In the rotating frame, the tidal potential allows for the spatially varying centrifugal force, with the tidal force $2q\Omega^{2}x\hat{x}$ exactly balanced by the Coriolis force $-2\Omega\hat{z}\times\mathbf{v}$ for circular orbits with $\mathbf{v}=\mathbf{v}_{0}$. For reference, Oort’s $A$ constant is related to $q$ by $A=q\Omega/2$, and the radial epicyclic frequency is given by $\kappa=\sqrt{2(2-q)}\Omega$. The simulations studied here fix $q=1$, corresponding to a flat rotation curve.
The key advantage of the shearing box setup from a numerical standpoint is the ability to construct self-consistent “shearing-periodic” boundary conditions in the plane, as illustrated in, e.g., Figure 1 of J. F. Hawley et al. (1995). The $y$ direction is purely periodic with $f(x,y,z)=f(x,y+L_{y},z)$ for any quantity $f$. However, upon crossing the box boundary in the ±$\pm x$ direction, the shear flow leads to an offset in $y$ of $\Delta y(t)=\mp q\Omega L_{x}(t-t_{n})$, where $t_{n}\equiv nL_{y}/(q\Omega L_{x})$ with $n\equiv\lfloor q\Omega tL_{x}/L_{y}\rceil$ is the “nearest periodic time” to $t$ and $\lfloor u\rceil$ denotes the closest integer to $u$. The $x$ boundary condition is thus $f(x,y,z)=f(x+L_{x},y+\Delta y(t),z)$ for all fluid variables other than $v_{y}$, for which instead (accounting for the shear across the box) $\delta v_{y}\equiv v_{y}+q\Omega x$, the azimuthal velocity relative to the circular orbit at radius $R_{0}+x$, satisfies this condition. These boundary conditions can be more succinctly expressed with the use of “sheared coordinates” $(x^{\prime},y^{\prime})$ given by
@@
×$\times e^{i[k_{x}^{\prime}(k_{x},k_{y},t)x^{\prime}(x)+k_{y}^{\prime}(k_{y})y^{\prime}(x,y,t)]},$ (6)
@@
The two particular TIGRESS-NCR simulations we study here are (i) the “R8-4 pc” model, with a nominal Galactocentric radius of $R_{0}=8\,$kpc and a spatial resolution of $\Delta x=4\,$pc, and (ii) the “LGR4-2 pc” model, with a nominal Galactocentric radius of $R_{0}=4\,$kpc and a spatial resolution of $\Delta x=2\,$pc. Table 1 summarizes the simulation parameters for each model. The simulations reach a statistical steady-state by $t\approx 200\,$Myr (see, e.g., Section 3.1 of C.-G. Kim et al. 2023). For the time interval $t\in[250\,\mathrm{Myr},450\,\mathrm{Myr}]$ and $t\in[250\,\mathrm{Myr},350\,\mathrm{Myr}]$ for the R8 and LGR4 simulations, respectively, the basic ISM properties and star formation rates (SFRs) have been analyzed in C.-G. Kim et al. (2023), and the radiation fields have been characterized in N. B. Linzer et al. (2024). Here, we characterize their surface density and potential fields every $\Delta t=1\,$Myr for the same intervals; we denote the temporal baseline with $t_{\mathrm{sim}}$ ($=200\,$Myr for R8 or 100100100\,100Myr for LGR4).
We choose these two simulations to span a range of gas surface densities and ISM conditions, representative of the gas structures in, e.g., nearby disk galaxies observed in the PHANGS survey (A. K. Leroy et al., 2021). The conditions in R8 are chosen to match Solar neighborhood conditions and are also similar to area-weighted averages for PHANGS, while conditions in LGR4 are similar to star-formation weighted averages from PHANGS (J. Sun et al., 2022). In principle, the analysis we present in Section 3 and Section 4 could be straightforwardly extended to simulations incorporating a strong spiral arm potential as in W.-T. Kim et al. (2020); we defer this generalization to future work.
@@
3 Surface Density
In this section, we will quantify the fluctuations in the surface density field. Note that here and throughout the remainder of this paper we will use the vector $\mathbf{x}$ to denote the planar variables only, i.e., $\mathbf{x}\equiv(x,y)$, and consider $z$-dependence separately.
Figure 2: A snapshot of the linear surface density fluctuation $\delta$ (left) and logarithmic surface density fluctuation $s$ (right) for the R8 simulation at $t=350\,$Myr. This is the same snapshot as in panel (c) of Figure 1.
@@ 3 Surface Density
Below, we will characterize the fields $\delta$ and $s$ by their spatial one-point (i.e., the pdfs of fluctuation values — see Section 3.1) and two-point (i.e., the fluctuation power spectra — see Section 3.2) statistics. We also describe the two-point spatio-temporal structure of the field in Section 3.3. Although these statistics are formally dependent on the grid cell size of the simulations used to study them, in Appendix A we demonstrate that our key results are converged with respect to spatial resolution by comparing measurements from our fiducial R8 simulation and a lower-resolution simulation with identical initial conditions. Note also that for the strongly nonlinear fluctuations (e.g., where $\delta\gtrsim 1$), the phases of different Fourier components of the field are correlated and information is therefore contained in higher order (beyond two-point) statistics; we discuss this caveat in Section 5.3.
@@
3.1 One-Point Spatial Surface Density Statistics
Figure 3: Histograms of the values of the logarithmic density fluctuation $s(\mathbf{x},t)$ measured at the simulation grid scale for the R8 (left, purple) and LGR4 (right, orange) simulations. The solid curves show the time-averaged histograms, while the shaded regions indicate ±$\pm 1\sigma$ variations about the values in each bin. A Gaussian with standard deviation $\sigma_{s}$ (corresponding to the measured standard deviation in the time-averaged values) and mean $\mu_{s}=-\sigma_{s}^{2}/2$ is overplotted as a black dashed curve.
In Figure 3 we show the area-weighted pdf $p(s)$ of the logarithmic surface density fluctuation $s(\mathbf{x},t)$ measured at the simulation grid scale for the R8 (left, purple) and LGR4 (right, orange) simulations. The dark curves in each panel show the time-averaged histogram across the entire temporal baseline we consider for each simulation, while the shaded region shows the ±$\pm 1\sigma$ variation about the time-averaged histogram. As anticipated from numerical simulations of isothermal, supersonic turbulence (e.g., E. Vazquez-Semadeni 1994; E. C. Ostriker et al. 2001), we find that the histograms near the peak are well-fit by Gaussian distributions [^3] So, the distribution of the surface density itself is log-normal.,
@@ 3.1 One-Point Spatial Surface Density Statistics
i.e., we weight each component $s_{\mathbf{k}}$ of the unsmoothed field with a window function $W^{(R)}$, and produce the smoothed field by summing up these components as in Section 2. The window functions are chosen to be real functions of $\mathbf{k}$ normalized so that $\int\mathrm{d}^{2}\mathbf{x}\,W^{(R)}(\mathbf{x})=1$ and $W^{(R)}(\mathbf{x}=0)=1/(\pi R^{2})$ for all $R$. This smoothing corresponds to the typical window convolution operation for a continuous and periodic field $Q$,
@@
We will show in Section 3.2 that the power spectrum $P_{s}$ can be approximately modeled as an isotropic power law in wavevectors $\mathbf{k}$, and so it is appropriate to use window functions that are isotropic in $\mathbf{k}$ as well. Since $\Delta x/L\ll 1$, we can approximate the sum over wavevectors as an integral; under the assumption of isotropy, we find
@@
Figure 5 shows the standard deviation measured directly from the smoothed field snapshots $s^{(R)}$ for each simulation (again with R8 in purple and LGR4 in orange) as a function of the smoothing scale $R$, with the time-average shown as a solid curve and the ±$\pm 1\sigma$ variation around the time-average shown as a shaded region. In the left panel we smooth the field using $W^{(R)}_{k}$, while in the right panel we use $W^{(R)}_{x}$. Both panels show similar amplitudes of $\sigma_{s}$ at each scale $R$, although the differences in the functional form of the window lead to slight deviations between them [^5] For instance, the measurement for $\sigma_{s}(R)$ derived from the sharp $k$-space window is less smooth and goes to zero at $R=L/\pi$ simply because it is more sensitive to the finite $k$-space resolution than the sharp real-space window, for which the weights assigned to each cell in $k$-space can take on values besides 00 or $1$.. For both simulations, we see that the widths of the shaded regions in Figure 5 are only weakly dependent on $R$, suggesting that feedback processes, shear, and turbulence contribute to time-variation across all scales similarly.
For comparison, in Figure 5 we also overplot dashed curves calculated using Section 3.1 (taking the reference scale $R_{\mathrm{ref}}$ to be the grid scale $\Delta x$ of each simulation) and the approximation Equation 18 with the simplified isotropic power-law model for $P_{s}$ described in Section 3.2 below. For the sharp $k$-space window $W^{(R)}_{k}$, this can be done analytically: substituting, we find
@@
×$\times\bigg{(}\bigg{(}\frac{\pi R}{L}\bigg{)}^{n_{s}-2}-\bigg{(}\frac{\pi\Delta x}{L}\bigg{)}^{n_{s}-2}\bigg{)}\bigg{]}^{1/2},$ (21)
@@
3.2 Two-Point Spatial Surface Density Statistics

### v2

@@ I Introduction
However, in reality, individual ISM substructures are observed to be far from spherically symmetric, often instead taking a more extended, filamentary form (e.g., J. D. Soler et al. 2022; A. Hacar et al. 2023 and references therein), and substantial fluctuations are observed in the Galaxy across a wide range of scales in atomic as well as molecular gas, and also in dust (e.g., Planck Collaboration et al., 2015). As resolution has improved, surveys of external galaxies in both the Local Group and to distances of tens of Mpc have also revealed extensive structure in both the gas and the dust (e.g., A. D. Bolatto et al., 2008; E. Rosolowsky et al., 2021; J. C. Lee et al., 2023). Thus, empirical characterization of ISM fluctuations now often take the form of multi-scale statistics rather than GMC (or clump) mass functions (see Section V.2).
In parallel with the development of higher resolution observations that enable more sophisticated empirical characterization of ISM structure, the advent of numerical hydrodynamics as a tool to investigate the ISM has led to vastly more realistic theoretical models. Different kinds of simulations provide different kinds of insight to ISM structure. Global galaxy and cosmological zoom simulations are able to capture important large-scale, long-lived features such as grand design spiral arms by solving N-body and fluid dynamics equations together, and including tidal gravity from nearby companion systems (e.g., T. Tepper-García et al. 2024). However, these simulations generally lack the spatial resolution and key ISM microphysical processes that would be needed to realistically model ISM fluctuations at scales below $\sim 100$ pc (with $\lesssim 10$ Myr lifetimes), since the computational cost would be prohibitive for galaxies larger than dwarfs (e.g., R. S. Somerville & R. Davé, 2015; T. Naab & J. P. Ostriker, 2017).
@@
In this work, we characterize ISM fluctuations that develop in the state-of-the-art TIGRESS-NCR simulations (C.-G. Kim et al., 2023). These simulations incorporate the effects of galactic rotation and magnetic fields while employing self-consistent star formation and feedback (ray-tracing radiation and supernovae) as well as nonequilibrium chemistry and cooling, with a simulation resolution of 242-4 pc. The simulations we analyze span both solar neighborhood and inner-galaxy conditions relevant in a variety of dynamical contexts. With an eye toward extragalactic observations of face-on disk systems (as in e.g., the PHANGS project, A. K. Leroy et al. 2021; J. C. Lee et al. 2023), we build a model for the ISM’s 2D surface density fluctuations, characterizing both their one-point and two-point statistics via probability density functions (pdfs) and power spectra. Additionally, we characterize the time dependence at each spatial frequency. For use in stellar dynamical applications, we also characterize the ISM’s 3D gravitational potential fluctuations.
The rest of this paper is organized as follows. First, in Section II, we describe in greater detail the TIGRESS-NCR simulations that we study throughout this work. In Section III and Section IV respectively, we characterize the ISM’s surface density fluctuations and connect them to the structure of the ISM’s volume density and gravitational potential. We discuss interpretations, implications, and caveats of our results, and connect with observations in Section V. Finally, we summarize in Section VI.
II Simulations
@@
The TIGRESS simulations employ the “shearing box” approximation (P. Goldreich & D. Lynden-Bell, 1965; J. F. Hawley et al., 1995) to model a locally Cartesian patch of the Galactic disk centered at a Galactocentric radius $R_{0}$, rotating at the local circular orbital frequency $\Omega$. The in-plane coordinates $x\equiv R-R_{0}$ and $y\equiv R_{0}(\varphi-\Omega t)$ represent the radial and azimuthal directions respectively, while the $z$ coordinate represents the distance above or below the midplane. The simulation domain extends a length $L_{x}=L_{y}\equiv L$ in each of the in-plane directions and $L_{z}$ in the vertical direction, with a cell size of $\Delta x$ in all dimensions, so that there are $N_{x}=N_{y}=L/\Delta x$ cells in each planar direction, and $N_{z}=L_{z}/\Delta x$ cells resolving the vertical direction. The background differential rotation of the galaxy is modeled as a linear shear flow, so that a circular orbit at a radial position $x$ has a velocity $\mathbf{v}_{0}=-q\Omega x\hat{y}$, where $q\equiv-(\mathrm{d}\ln\Omega/\mathrm{d}\ln R)_{R_{0}}$ parametrizes the shear. In the rotating frame, the tidal potential allows for the spatially varying centrifugal force, with the tidal force $2q\Omega^{2}x\hat{x}$ exactly balanced by the Coriolis force $2\Omega\hat{z}\times\mathbf{v}$ for circular orbits with $\mathbf{v}=\mathbf{v}_{0}$. For reference, Oort’s $A$ constant is related to $q$ by $A=q\Omega/2$, and the radial epicyclic frequency is given by $\kappa=\sqrt{2(2-q)}\Omega$. The simulations studied here fix $q=1$, corresponding to a flat rotation curve.
The key advantage of the shearing box setup from a numerical standpoint is the ability to construct self-consistent “shearing-periodic” boundary conditions in the plane, as illustrated in, e.g., Figure 1 of J. F. Hawley et al. (1995). The $y$ direction is purely periodic with $f(x,y,z)=f(x,y+L_{y},z)$ for any quantity $f$. However, upon crossing the box boundary in the $\pm x$ direction, the shear flow leads to an offset in $y$ of $\Delta y(t)=\mp q\Omega L_{x}(t-t_{n})$, where $t_{n}\equiv nL_{y}/(q\Omega L_{x})$ with $n\equiv\lfloor q\Omega tL_{x}/L_{y}\rceil$ is the “nearest periodic time” to $t$ and $\lfloor u\rceil$ denotes the closest integer to $u$. The $x$ boundary condition is thus $f(x,y,z)=f(x+L_{x},y+\Delta y(t),z)$ for all fluid variables other than $v_{y}$, for which instead (accounting for the shear across the box) $\delta v_{y}\equiv v_{y}+q\Omega x$, the azimuthal velocity relative to the circular orbit at radius $R_{0}+x$, satisfies this condition. These boundary conditions can be more succinctly expressed with the use of “sheared coordinates” $(x^{\prime},y^{\prime})$ given by
@@
$\times e^{i[k_{x}^{\prime}(k_{x},k_{y},t)x^{\prime}(x)+k_{y}^{\prime}(k_{y})y^{\prime}(x,y,t)]},$ (6)
@@
The two particular TIGRESS-NCR simulations we study here are (i) the “R8-4 pc” model, with a nominal Galactocentric radius of $R_{0}=8\,$kpc and a spatial resolution of $\Delta x=4\,$pc, and (ii) the “LGR4-2 pc” model, with a nominal Galactocentric radius of $R_{0}=4\,$kpc and a spatial resolution of $\Delta x=2\,$pc. Table 1 summarizes the simulation parameters for each model. The simulations reach a statistical steady-state by $t\approx 200\,$Myr (see, e.g., Section 3.1 of C.-G. Kim et al. 2023). For the time interval $t\in[250\,\mathrm{Myr},450\,\mathrm{Myr}]$ and $t\in[250\,\mathrm{Myr},350\,\mathrm{Myr}]$ for the R8 and LGR4 simulations, respectively, the basic ISM properties and star formation rates (SFRs) have been analyzed in C.-G. Kim et al. (2023), and the radiation fields have been characterized in N. B. Linzer et al. (2024). Here, we characterize their surface density and potential fields every $\Delta t=1\,$Myr for the same intervals; we denote the temporal baseline with $t_{\mathrm{sim}}$ (=200=200\,Myr for R8 or 100100\,Myr for LGR4).
We choose these two simulations to span a range of gas surface densities and ISM conditions, representative of the gas structures in, e.g., nearby disk galaxies observed in the PHANGS survey (A. K. Leroy et al., 2021). Note that these simulations are not intended to model different regions within a single galaxy: the conditions in R8 are chosen to match solar neighborhood conditions and are also similar to area-weighted averages from PHANGS, while conditions in LGR4 are similar to star-formation-weighted averages from PHANGS (see Table 1 of J. Sun et al., 2022). In principle, the analysis we present in Section III and Section IV could be straightforwardly extended to simulations incorporating a strong spiral arm potential as in W.-T. Kim et al. (2020); we defer this generalization to future work.
@@
III Surface Density
In this section, we will quantify the fluctuations in the surface density field, calculated by integrating the volume density over the vertical extent of the box. Note that here and throughout the remainder of this paper we will use the vector $\mathbf{x}$ to denote the 2D planar variables only, i.e., $\mathbf{x}\equiv(x,y)$, and consider $z$-dependence explicitly when needed.
Figure 2: A snapshot of the linear surface density fluctuation $\delta$ (left) and logarithmic surface density fluctuation $s$ (right) for the R8 simulation at t=350t=350\,Myr. This is the same snapshot as in panel (c) of Figure 1.
@@ III Surface Density
Below, we will characterize the fields $\delta$ and $s$ by their spatial one-point (i.e., the pdfs of fluctuation values—see Section III.1) and two-point (i.e., the fluctuation power spectra—see Section III.2) statistics. We also describe the two-point spatio-temporal structure of the field in Section III.3. Although these statistics are formally dependent on the grid cell size of the simulations used to study them, in Appendix A we demonstrate that our key results are converged with respect to spatial resolution by comparing measurements from our fiducial R8 simulation and a lower-resolution simulation with identical initial conditions. Note also that for the strongly nonlinear fluctuations (e.g., where $\delta\gtrsim 1$), the phases of different Fourier components of the field are correlated and information is therefore contained in higher order (beyond two-point) statistics; we discuss this caveat in Section V.3.
@@
III.1 One-Point Spatial Surface Density Statistics
Figure 3: Histograms of the values of the logarithmic density fluctuation $s(\mathbf{x},t)$ measured at the simulation grid scale for the R8 (left, purple) and LGR4 (right, orange) simulations. The solid curves show the time-averaged histograms, while the shaded regions indicate $\pm 1\sigma$ variations about the values in each bin. A Gaussian with standard deviation $\sigma_{s}$ (corresponding to the measured standard deviation in the time-averaged values) and mean $\mu_{s}=-\sigma_{s}^{2}/2$ is overplotted as a black dashed curve.
In Figure 3 we show the area-weighted pdf $p(s)$ of the logarithmic surface density fluctuation $s(\mathbf{x},t)$ measured at the simulation grid scale for the R8 (left, purple) and LGR4 (right, orange) simulations. The dark curves in each panel show the time-averaged histogram across the entire temporal baseline we consider for each simulation, while the shaded region shows the $\pm 1\sigma$ variation about the time-averaged histogram. As anticipated from numerical simulations of isothermal, supersonic turbulence (e.g., E. Vazquez-Semadeni 1994; E. C. Ostriker et al. 2001), we find that the histograms near the peak are well-fit by Gaussian distributions [^3] So, the distribution of the surface density itself is log-normal.,
@@ III.1 One-Point Spatial Surface Density Statistics
i.e., we weight each component $s_{\mathbf{k}}$ of the unsmoothed field with a window function $W^{(R)}$, and produce the smoothed field by summing up these components as in Section II. The window functions are chosen to be real functions of $\mathbf{k}$ normalized so that $\int\mathrm{d}^{2}\mathbf{x}\,W^{(R)}(\mathbf{x})=1$ and $W^{(R)}(\mathbf{x}=0)=1/(\pi R^{2})$ for all $R$. This smoothing corresponds to the typical window convolution operation for a continuous and periodic field $Q$,
@@
We will show in Section III.2 that the power spectrum $P_{s}$ can be approximately modeled as an isotropic power law in wavevectors $\mathbf{k}$, and so it is appropriate to use window functions that are isotropic in $\mathbf{k}$ as well. Since $\Delta x/L\ll 1$, we can approximate the sum over wavevectors as an integral; under the assumption of isotropy, we find
@@
Figure 5 shows the standard deviation measured directly from the smoothed field snapshots $s^{(R)}$ for each simulation (again with R8 in purple and LGR4 in orange) as a function of the smoothing scale $R$, with the time-average shown as a solid curve and the $\pm 1\sigma$ variation around the time-average shown as a shaded region. In the left panel we smooth the field using $W^{(R)}_{k}$, while in the right panel we use $W^{(R)}_{x}$. Both panels show similar amplitudes of $\sigma_{s}$ at each scale $R$, although the differences in the functional form of the window lead to slight deviations between them [^5] For instance, the measurement for $\sigma_{s}(R)$ derived from the sharp $k$-space window is less smooth and goes to zero at $R=L/\pi$ simply because it is more sensitive to the finite $k$-space resolution than the sharp real-space window, for which the weights assigned to each cell in $k$-space can take on values besides 0 or $1$.. For both simulations, we see that the widths of the shaded regions in Figure 5 are only weakly dependent on $R$, suggesting that feedback processes, shear, and turbulence contribute to time-variation across all scales similarly.
For comparison, in Figure 5 we also overplot dashed curves calculated using Section III.1 (taking the reference scale $R_{\mathrm{ref}}$ to be the grid scale $\Delta x$ of each simulation) and the approximation Equation 18 with the simplified isotropic power-law model for $P_{s}$ described in Section III.2 below. For the sharp $k$-space window $W^{(R)}_{k}$, this can be done analytically: substituting, we find
@@
$\times\bigg(\bigg(\frac{\pi R}{L}\bigg)^{n_{s}-2}-\bigg(\frac{\pi\Delta x}{L}\bigg)^{n_{s}-2}\bigg)\bigg]^{1/2},$ (21)
@@
III.2 Two-Point Spatial Surface Density Statistics